OpenStudy (anonymous):

find the total area of cylinder whose radius is 2 Cm and height is 4 cm

OpenStudy (anonymous):

TSA= 2 Pi r^2 + 2 pirh =2 rpi (r+h) =circumference (radius + height TSA= ((pi x 2(2)) (2+4) =4pi (6) =24pi =75.39822369 cm^2